PLS HELP ME WITH MY PROBLEM!!
QUADRATIC EQUATION BY COMPLETING SQUARE
12X^2+11X+2=0

Answer by stanbon(75887)

QUADRATIC EQUATION BY COMPLETING SQUARE
12X^2+11X+2=0
----------------------
Reform the equation:
12(x^2 + (11/12)x + ?) = -2 +12*?
---
12(x^2 + (11/12)x +(11/24)^2) = -2 + 12(11/24)^2
---
12(x+(11/24))^2 = -2 + 121/48
12(x+(11/24))^2 = (-96/48)+(121/48)
---
12(x+(11/24))^2 = 25/48
---
(x+(11/24))^2 = 25/576
------
x = (5/24)-(11/24) or x = (-5/24)-(11/24)
---
x = -6/24 or x = -16/24
x = -1/4 or x = -2/3
=================================
Cheers,
Stan H.